OpenStudy (loser66):

Glenn invested $18,000 at 5% interest compounded annually. How much interest will Glenn earn in 3 years? A. $2837.25 B. $19,826.24 C. $20,837.25 D. $2340.23

OpenStudy (loser66):

@Michele_Laino

OpenStudy (anonymous):

Umm>

iYuko (iyuko):

Do you know the interest formula?

OpenStudy (loser66):

nope

OpenStudy (loser66):

I would like to know the logic how to get the answer, not the formula.

OpenStudy (anonymous):

I found this from another open study person: On how they got A simple interest, I = Principal,p * rate as a %, r/100* time ,t in years I = P RT/100 p = 18000; r = 5% = 5/100; t = 3 years plug the numbers compound interest I =P (1 + r)^t - P where r is the rate as a % and t = number of compounding periods

OpenStudy (anonymous):

c

iYuko (iyuko):

I=PRT

OpenStudy (loser66):

I got c also, but it is wrong and the correct one is d. How??? no formula, logic, please

OpenStudy (anonymous):

I can copy and paste the review question on how it was wrong and why??

OpenStudy (loser66):

@catterang to me, I believe things I see the proof, not the answer.

OpenStudy (anonymous):

This is from k12..

OpenStudy (anonymous):

They explain stuff

OpenStudy (anonymous):

This is what they said: 5. Glenn invested $18,000 at 5% interest compounded annually. How much interest will Glenn earn in 3 years? A. $2837.25 B. $19,826.24 C. $20,837.25 D. $2340.23 Incorrect A. Use the formula to compute A, the total amount of money Glenn would have after 3 years: A = P(1 + r)t. Substitute the values you know: P = 18,000; r = 0.05; t = 3: A = P(1 + r)t A = 18,000(1 + 0.05)3 Substitute. A = 18,000(1.05)3 Add values in parentheses. A = 18,000(1.157625) Compute the power. A = 20,837.25 Multiply. Since $20,837.25 is the total amount of money, subtract the original principal of $18,000 to find the amount of interest paid: $20,837.25 – $18,000 = $2837.25 of interest C. Use the formula to compute A, the total amount of money Glenn would have after 3 years: A = P(1 + r)t. Substitute the values you know: P = 18,000; r = 0.05; t = 3: A = P(1 + r)t A = 18,000(1 + 0.05)3 Substitute. A = 18,000(1.05)3 Add values in parentheses. A = 18,000(1.157625) Compute the power. A = 20,837.25 Multiply. Since $20,837.25 is the total amount of money, subtract the original principal of $18,000 to find the amount of interest paid: $20,837.25 – $18,000 = $2837.25 of interest

OpenStudy (loser66):

@ikram002p lalala, my Master is here. Help me, friend

OpenStudy (anonymous):

Ya answer is A

OpenStudy (loser66):

oooooooooooooooh, got it. they ask about the INTEREST, not the TOTAL AMOUNT after 3 years. ha!!!! ENGLISH mistake

OpenStudy (anonymous):

Here is the way you compute it (18,000(1+.05)^3)-18,000

OpenStudy (anonymous):

yeaaa lol :D

OpenStudy (anonymous):

Yes

OpenStudy (loser66):

Thanks you guys. Thanks for making it clear to me. hehehe..... if not, I can't sleep tonight

OpenStudy (anonymous):

hahahahah

OpenStudy (anonymous):

Loser can you check my problem?